The average weight of 25 chocolate bars selected from a normally distributed population is 200 g with a standard deviation of 10 g. Find the interval estimate using 98% confidence level.


1
Expert's answer
2022-05-24T09:49:53-0400

The critical value for "\alpha = 0.02" is "z_c = z_{1-\alpha/2} = 2.3263."

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{x}+z_c\times\dfrac{\sigma}{\sqrt{n}})"

"=(200-2.3263\times\dfrac{10}{\sqrt{25}}, 200+2.3263\times\dfrac{10}{\sqrt{25}})"

"=(195.3474, 204.6526)"

Therefore, based on the data provided, the 98% confidence interval for the population mean is "195.3474 < \mu < 204.6526," which indicates that we are 98% confident that the true population mean "\mu" is contained by the interval "(195.3474, 204.6526)."